40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 27 02:09:14 2015, ericq, Update, LSC, AUX X Laser Current Changed 
    Reply  Fri Sep 18 11:59:08 2015, ericq, Update, LSC, AUX X Laser Current Reverted 
Message ID: 11619     Entry time: Fri Sep 18 11:59:08 2015     In reply to: 11533
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX X Laser Current Reverted 

Once again, the transmitted X green beam was showing enormous intensity fluctuations (50x higher than normal). Last month, I reduced the AUX X laser current from 2.0A to 1.9A, which I thought had fixed it somehow.

However, when I sent to the end to check it out today, I found the SR560 which is there to amplify the green PDH error signal before being sent to the AA board was overloading. Not so surprising, since the error signal was similarly noisy as the transmitted light. 

I turned the SR560 gain down, and, after relocking, the transmitted light was stable. I've turned the AUX X laser current back up to 2.0A, it's previous nominal value, and the green transmitted light is still stable. 

I'm a little mystified that the 560 could intefere with the loop, since it is not in the feedback path. Could it be that when it is overloading, it sends garbage backwards out of the inputs? But even then, its input is not connected to the real error point, but the buffered monitor port. Could it be interfering via the power line?

Before, I had hesitated adding gain to the PDH board's monitor point for DAQ purposes, because the motivation of the port is to provide a 1:1 version of the real error signal, and I didn't want to add gain to the AA board, because we normally don't have gain in those boards, and I didn't want to surprise future people. The SR560 was meant to be temporary, but as often happens, it was forgotten. Now, I think I will add gain to the error monitor buffer stage of the PDH boards. 

ELOG V3.1.3-